1. An arrow accelerates from 0 to 40 m/s in the 0.5 s it is in contact with the bow string. What is the average acceleration?

2. A coastal force of 20 N acts on a mass of (a) 2 kg, (b) 4 kg, (c) 6 kg. What are the resulting acceleration?

3. A 0.5 kg wrench is dropped from a height of 10 m. What is the momentum just before it strikes the floor?
1
Expert's answer
2020-01-30T09:42:30-0500

1.The average acceleration


"a=\frac{\Delta v}{\Delta t}=\frac{40-0}{0.5}=80\:\rm m/s^2"

2. The Newton's second law states


"a=\frac{F}{m}"

Hence

(a) "a=\frac{20}{2}=10\:\rm m/s^2"

(b) "a=\frac{20}{4}=5\:\rm m/s^2"

(c) "a=\frac{20}{6}=3.3\:\rm m/s^2"


3. The momentum


"p=mv=m\sqrt{2gh}""=0.5\times \sqrt{2\times 9.8\times 10}=7\:\rm kg\cdot m/s"